The question is: Use Elimination to solve:
7x-4y=29
3x+5y=-1

Answer by stanbon(75887)

You can choose to eliminate the x's or the y's.
Let's say you want to eliminate the y's.
Multiply the 1st equation by 5 to get:
1) 35x-20y=145
Multiply the 2nd equation by 4 to get:
2) 12x+20y=-4
Now, subtract 2) from 1) to get:
23x+0y=149
Now you can solve this for "x"
Then substitute the value you get for "x"
into one of the original equations to solve
for "y".
Hope this helps.
Cheers,
Stan H.